whitemonsterbunny17 (whitemonsterbunny17):

Help with significant figures? Thanks!

whitemonsterbunny17 (whitemonsterbunny17):

15.75 m x 8.45 m

OpenStudy (preetha):

White monster, can you please type using the usual script!

whitemonsterbunny17 (whitemonsterbunny17):

Sure.

OpenStudy (anonymous):

Multiply normally but look at the units, m*m = ?

whitemonsterbunny17 (whitemonsterbunny17):

Yeah, what do I do with those?

OpenStudy (anonymous):

\[15.75m*8.45m = 133.1m^2\]

whitemonsterbunny17 (whitemonsterbunny17):

Oh, so I just add an exponent of 2 to them?

OpenStudy (anonymous):

Think of it as m^1 and treat it as if they were exponents :P

OpenStudy (anonymous):

m^1*m^1 = m^(1+1)

whitemonsterbunny17 (whitemonsterbunny17):

And that's only for if they are the same unit, right?

OpenStudy (anonymous):

yup

whitemonsterbunny17 (whitemonsterbunny17):

Okay, cool. So it should be: 133.1 m^2

OpenStudy (anonymous):

yup

whitemonsterbunny17 (whitemonsterbunny17):

YAY!! \(\Large\bf\rlap{\color{red}{\bigstar~T}\color{gold}{H}\color{greenyellow}{A}\color{lime}{N}\color{cyan}{K~}\color{blue}{Y}\hspace{0 pt}\color{purple}{O}\color{magenta}{U}\color{orchid}{!~\bigstar}}{\color{black}{\; \bigstar~THANK~YOU!~\bigstar}}\)

OpenStudy (anonymous):

Np ^.^